# LLM-guided Chemical Process Optimization with a Multi-Agent Approach

### v1

@@ 4 Conclusions
The framework’s primary innovation lies in autonomous constraint generation combined with collaborative multi-agent optimization. By leveraging embedded domain knowledge to generate engineering-appropriate operating bounds for unit operations, the system enables optimization in scenarios where traditional methods cannot be directly applied. Combined with flexibility in handling non-differentiable objective functions, this capability enables optimization where traditional gradient-based methods may encounter difficulties. Validation on the HDA process demonstrated competitive performance against IPOPT and grid search across multiple objectives, with significant computational efficiency gains achieving convergence in under 20 minutes compared to over 10 hours for grid search. Notably, the framework’s reasoning capabilities distinguish it from conventional optimizers, with agents demonstrating sophisticated process understanding through explicit trade-off identification and domain-informed parameter selection strategies.
@@ 5 Code Availability
[…]
{suppinfo}

### v2

@@ 4 Conclusions
The framework’s primary innovation lies in autonomous constraint generation combined with collaborative multi-agent optimization. By leveraging embedded domain knowledge to infer engineering-appropriate operating bounds for unit operations, the system enables optimization in scenarios where traditional methods cannot be directly applied. The framework’s flexibility in handling black-box and non-differentiable objective functions, combined with its ability to provide natural language explanations for parameter choices, addresses key limitations of conventional optimizers. Validation on the HDA process demonstrated competitive performance against IPOPT and grid search across multiple objectives, with significant computational efficiency gains achieving convergence in under 20 minutes compared to over 10 hours for grid search. Beyond computational metrics, the framework’s reasoning capabilities provide interpretable optimization trajectories, with agents explicitly articulating process trade-offs and domain-informed parameter selection strategies that enhance engineer trust and facilitate safety validation.
@@
https://github.com/tongzeng24/ProcessAgent
Supporting Information